OpenStudy (anonymous):

Why are the solutions to 4sin(x)cos(x)=1 (π/12),(5π/12),(13π/12) and (17π/12)? I understand the (π/12), but there are unsymmetrical gaps in between the answers to the equation.

OpenStudy (loser66):

show me how you get those answer, please

OpenStudy (anonymous):

I got (π/12) myself, the rest are from the answers from the back of the book. (There's no point in writing them down if you don't know how you got them.)

OpenStudy (anonymous):

4sin(x)cos(x)=1 2sin(2x)=1 *(1/2) *(1/2) sin(2x)=(1/2) sinx(1/2) at (π/6) (π/6)/(1/2) = (π/12)

OpenStudy (anonymous):

maybe I did that wrong, I don't really know

OpenStudy (loser66):

you are very good at this. ok. now I set 2x = \(\theta\) got me?

OpenStudy (anonymous):

*sinx=(1/2) at (π/6)

OpenStudy (anonymous):

I am aware of adding 90 degrees (or something) to get every answer from 0 to 2π, but the answer don't follow that.

OpenStudy (loser66):

so, for 1/2 you have 2 values of \(\theta\) not just 1. I show you about \(\huge \theta\) \(sin\theta = \dfrac{1}{2}\) \(\rightarrow \theta =\dfrac{\pi}{6}~~and~~ \theta = \dfrac{5\pi}{6}\)